OpenStudy (anonymous):

Write the standard equation for the circle that passes through the points: (0, 0) (6, 0) (0, - 8)

OpenStudy (anonymous):

For my first equation I got f=0 because \[0^2 + 0^2 + 0d + 0e +f =0\] Then for my second equation I got \[-36=6d+f\] For my third equation I got \[-64=-8e\] The problem I have now is that I have to subtract my second equation from my first equation, but the first equation isn't really an equation I don't think. It's just zero. So would I just turn \[-36 = 6d + f ...into... 36 = -6d - f\]

OpenStudy (nurali):

Write the standard equation for the circle that passes through the points: (0, 0) (6, 0) (0, - 8) Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k) being the (x,y) coordinates of the center, r=radius. Solving for h, k and r, using given points .. Equation for point (0,0): (0-h)^2+(0-k)^2=r^2 eq1) h^2+k^2=r^2 .. Equation for point (6,0): (6-h)^2+(0-k)^2=r^2 eq2) (6-h)^2+k^2=r^2 .. Equation for point (0,-8): (0-h)^2+(-8-k)^2=r^2 eq3) h^2+(-8-k)^2=r^2 .. eq1) h^2+k^2=r^2 eq2) (6-h)^2+k^2=r^2 subtract eq2 from eq1 eliminating k^2 and r^2 h^2-(6-h)^2=0 h^2-(36-12h+h^2)=0 h^2-36+12h-h^2=0 12h=36 h=3 .. eq1) h^2+k^2=r^2 eq3) h^2+(-8-k)^2=r^2 subtract eq3 from eq1 eliminating h^2 and r^2 k^2-(-8-k)^2=0 k^2-(64+16k+k^2)=0 k^2-64-16k-k^2=0 -16k=64 k=-4 .. eq1) h^2+k^2=r^2 3^2+(-4)^2=r^2 9+16=r^2 r^2=25 .. Equation of given circle: (x-3)^2+(y+4)^2=25

OpenStudy (anonymous):

Well I was given the equation \[x^2 + y^2 + dx + ey + f = 0\] to use

OpenStudy (anonymous):

It's the same equation, just distribute it out. x-3 squared is x^2 -6x + 9 y+4 squared is y^2 + 8y + 16 All together, in your form, the equation would be: x^2 + y^2 -6x + 8y + 9 + 16 - 25 = 0

OpenStudy (anonymous):

That would be my final equation?

OpenStudy (anonymous):

simply the constants, which in this case would be zero. If your ever unsure, graph the points. From a quick sketch, you can find the radius and center. From there, you can use the equation Nurali used, then just work out the algebra like I did to find the other form. Try it now on some scratch paper to see if you know what I'm saying. It should take you like a minute tops

OpenStudy (anonymous):

Oh okay. Yeah graphing always helps. Thanks!!